# Information Discovery in e-Commerce

### v1

@@ 7.2.1 Introduction to dialogue systems
Generation-based dialogue systems aim to generate natural-sounding replies automatically to exchange information (e.g., knowledge, sentiments, etc.) and complete a variety of specific tasks in a conversation interaction process [Young2013POMDP, shawar2007]. End-to-end textual generation models [shang, vinyals2015neural, sordoni2015, li2016a, li2016b, serban2016building] have proved capable in multiple dialogue systems applications with promising performance. Most end-to-end neural generation models apply an encoder-decoder architecture based on a recurrent neural network, which directly maps an input context to the output response. Several approaches have been proposed to softly model language patterns, such as word alignment and repeating into sequence-to-sequence structure [bahdanau2015neural, gu2016incorporating, serban2017multiresolution, cao2017]. gu2016incorporating proposed a copy mechanism to consider additional copying probabilities for contextual words in forum conversations. serban2017multiresolution decoded coarse tokens before generating the complete response. Variational neural networks perform efficient inference and learning in models with directed probabilities on a large-scale dataset [Kingma2014Auto, Kingma2014Adam]. cao2017 tackled the boring output issue of deterministic dialogue models by introducing a latent variable model for a one-shot dialogue response. [hvred] proposed HVRED to utilize the latent variable at the sub-sequence level in a hierarchical setting, whereas ziwwsdm1 added a hierarchical structure and a variational memory module into a neural encoder-decoder network.
@@
Figure 7.5: An example e-commerce dialogue in the JDDC 1.0 dataset. Image source: [chen2019jddc].
@@
In e-commerce platforms, after-sale customer service is the main application scenario for dialogue systems. In contrast with existing dialogue systems, e-commerce dialogues need to address three targets: (i) task completion, such as changing the order address, providing the receipt, and returning the order; (ii) knowledge-based response selection and generation, such as checking the status of the delivery, answering the request about the refund period; and (iii) empathetic response generation, such as satisfying the consumers’ request and replying to consumers’ complaints. JDDC datasets are collected from JD.com. one of the largest e-commerce platforms in China. Fig. 7.5 shows a typical example dialogue from JDDC 1.0 dataset. We can see that the blue text shows the target completion task, the red text indicates the knowledge-based response generation task, and the purple text shows the empathetic response generation task. Another characteristic of e-commerce dialogue systems is the phenomenon of multiple modalities. It is observed that text and images are often used in customer service dialogues [zhao2021jddc, yuan2022mcic].
@@
Fig. 7.6 shows another example from the JDDC 2.0 dataset to demonstrate the multi-modality characteristic of e-commerce. In this figure, $3$ dialogue segments show images that are used to distinguish different product models for the same brand or used for identifying the location and cause of product failures. According to the above $3$ targets in e-commerce dialogues, we detail e-commerce dialogue systems through $3$ angles: task-oriented dialogue systems, knowledge-grounded dialogue systems, and empathetic dialogue systems in Sections 7.2.27.2.3, and 7.2.4, respectively.
@@ 7.2.2 Task-oriented dialogue systems
The methods of task-oriented dialogue systems can be divided into pipeline methods and end-to-end methods. As shown in Fig. 7.4, the pipeline of task-oriented dialogue systems can be divided into natural language understanding (NLU), dialogue state tracking (DST), and dialogue policy learning (DPL), and natural language generation (NLG) [Chen2017A]. For each stage, there are a bunch of pipeline-based approaches have been proposed. Despite a lot of domain-specific handcrafting in traditional task-oriented dialogue systems, which are difficult to adapt to new domains [bordes2016learning]. Recently, end-to-end neural network solutions have been widely applied to the task [bordes2016learning, zhao2016towards, wen2017network, jin2018explicit, gao2018neural, madotto2018mem2seq].
@@
In NLU, the dialogue system maps the input utterance into semantic slots. These semantic slots are pre-defined in different application scenarios. NLU includes two challenging problems: intent detection and slot filling. Intent detection methods for language understanding are performed to detect the user’s intent [deng2012use, tur2012towards]. Deep neural networks have also been applied to detect the user’s intent. huang2013learning use convolutional neural networks (CNN) to detect the user’s intent. A similar CNN-based classification framework also resembled [shen2014learning]. Slot filling is usually set as a sequence labeling problem, where words are assigned with semantic labels [Chen2017A]. Deep belief networks have been successfully applied to address the filling problem [deng2012use, deoras2013deep]. Subsequently, recurrent neural networks are verified as effective in slot filling [mesnil2013investigation, sarikaya2011deep, yao2013recurrent, yao2014spoken]. Unlike other NLU approaches, liang2020moss jointly formulated intent detection and slot filling as a sequence generation problem. rastogi2020towards provided a schema-guided paradigm for NLU. Recently, pre-trained language models have been applied to enhance NLC in task-oriented dialogue systems. wu2020tod proposed a self-supervised language model trained on multiple task-oriented dialogue system benchmarks. The authors verified that their proposed model achieved the best in-scope and out-of-scope performances in NLU. zhang2021effectiveness proposed a pre-trained model for few-shot NLU by fine-tuning BERT on a small set of labeled utterances. 10.1145/3477495.3532069 explored tree-induced semi-supervised contrastive pre-training for NLU in task-oriented dialogue systems. The authors improved the NLU performance by injecting structural-semantic information to enhance the representation of dialogues. To explore more knowledge from long sequences in dialogue context, zhong2022dialoglm proposed a window-based pre-trained model for NLU based on the sequence-to-sequence model architecture.
@@
In a conversation, a dialogue state refers to a full and temporal representation of each participant’s intention [goddeau1996form]. In task-oriented dialogue generation, dynamically tracking dialogue states is the key to generating coherent and context-sensitive responses. In DST, we use a dialogue state $H_{t}$ to denote the representation of the dialogue till time $t$. Traditional approaches to DST focus on searching hand-crafted rules to select the most likely results [wang2013simple, young2010hidden, williams2012belief], where the dialogue state tracking is transferred to a slot filling problem. Such a kind of slot filling problem was also addressed by approaches using conditional random fields [lee2013structured, lee2013recipe, ren2013dialog] and maximum entropy [williams2013multi]. However, relying on the most likely results from an NLU module [Perez17], these rule-based systems hardly model uncertainty, which is prone to frequent errors [williams2014web, Perez17].
@@
Unlike rule-based state tracking methods, young2010hidden proposed a distributional dialogue state for statistical dialog systems and maintained a distribution over multiple hypotheses facing noisy conditions and ambiguity. Neural networks have been successfully applied to dialogue state tracking [henderson2013deep, mrkvsic2015multi, mrkvsic2017neural]. In task-oriented dialogue systems, end-to-end neural networks are employed for tracking dialogue states via interacting with an external knowledge base [wen2017network, eric2017key, bordes2016learning, williamszweig2017Long]. wen2017network divided the training procedure into two phases: the dialogue state tracker training and the complete model training. mrkvsic2017neural proposed a dialogue state tracker based on word embedding similarities. eric2017key implicitly modeled a dialogue state through an attention-based retrieval mechanism to reason over a key-value representation of the underlying knowledge base. bordes2016learning tracked the dialogue context in a memory module and repeatedly searched this context to select an adequate system response. Instead of employing symbolic knowledge queries, dhingra-EtAl:2017:Long1 proposed an induced “soft” posterior distribution over the knowledge base to search for matching entities. williamszweig2017Long combined an RNN with domain-specific knowledge encoded as software and system action templates. The copying mechanism is verified effective as generative dialogue state tracking. lei2018sequicity proposed an extendable framework to track dialogue states with a text span, including the constraints for a knowledge base query. The limited amount of labeled data is a severe challenge for DST. jin2018explicit proposed a semi-supervised way to integrate a copy procedure with the dialogue state tracking.
@@
While early studies on DST methods focused on a single-domain scenario, recent studies have turned their attention to multi-domain DST with the release of a multi-domain DST benchmark dataset, MultiWoZ [budzianowski2018multiwoz]. ramadan2018large jointly identified the domain and tracks the belief states corresponding to that domain to address the multi-domain DST problem. zhou2019multi formulated multi-domain DST as a question-answering task and used reading comprehension techniques to generate the answers. Similarly, gao2019dialog also formulated DST as a reading comprehension task and proposed an attention-based neural network to find the state answer as a span over tokens. DSTC challenges have provided a series of popular experimentation frameworks and dialog datasets collected through human-machine interactions for benchmarking [henderson2014second, henderson2014third, williams2014dialog, williams2016dialog].
@@
In real-world scenarios, it is often not practical to enumerate all possible slot value pairs and perform scoring from a large, dynamically changing knowledge base [xu2018end]. wu2019transferable proposed a method to generate dialogue states from utterances using a copy mechanism, where tracking knowledge across domains is shared. To alleviate the data sparsity in DST, yin2020dialog proposed a reinforced data augmentation framework to increase both the amount and diversity of the training data. chen2020schema incorporated slot relations and model slot interactions in multi-domain dialogue state tracking to enhance the slot interrelation between disciplines. feng2022dynamic extended this method by dynamically updating slot relations in the schema graph. heck2020trippy maintained two memories in DST: one for system inform slots and one for the previously seen slots. li-etal-2021-generation proposed the generation and extraction combined method with hierarchical ontology integration for DST. To tackle the understanding of enormous ellipsis and reference expressions in open vocabulary-based methods, ouyang2020dialogue proposed a copy-augmented encoder-decoder model by connecting the target slot and its source slot explicitly. liao2021multi formulated multi-domain DST as a recursive inference mechanism to improve the generation performance. Most existing DST models are usually trained offline, which requires a fixed dataset prepared in advance. Given a new domain in multi-domain DST, campagna2020zero proposed a zero-shot transfer learning method to handle new domains without incurring the high cost of data acquisition. Recently, the granularity of dialogue history has been proposed to mitigate the sparseness in DST [yang2021comprehensive]. guo2022beyond proposed a multi-perspective dialogue collaborative selector module to dynamically select the granularity of dialogue history in DST.
@@
Pre-trained language models have been verified as effective in dialogue state tracking. lin2021zero applied a pre-trained model for DST to exploit external knowledge from reading comprehension data. Similarly, zhong2022dialoglm verified document summarization can provide helpful signals to improve DST. liu2021domain introduced domain-lifelong learning into DST. The authors proposed a knowledge preservation network that includes a multi-prototype enhanced retrospection component and a multi-strategy knowledge distillation component. lin2021leveraging successfully applied T5 to improve zero-shot cross-domain DST. lee2021dialogue proposed a solution for multi-domain DST by prompting knowledge from a large-scale pre-trained language model. lin2021knowledge proposed a hybrid method to integrate GPT-2 with graph attention networks to enhance the DST performance. To mitigate the problem of incorrect in DST, wang2022luna proposed a BERT-based method by explicitly aligning each slot with its most relevant utterance.
@@
Scalability, robustness, and efficiency in DST have also been addressed recently. lei2018sequicity proposed a two-stage copy-aware network demonstrating good scalability. ren2019scalable considered the DST task as a sequence generation problem and proposed a scalable hierarchical encoder-decoder neural network with constant inference time complexity. kumar2020ma extended to improve the encoding of dialog context and slot semantics for DST to robustly capture critical dependencies between slots and the conversation history. kim2020efficient focused on an open vocabulary-based setting and considered the dialogue state as a memory that can be selectively overwritten to improve the efficiency in multi-domain DST. zhu2020efficient proposed an efficient multi-domain dialogue state tracker by jointly encoding the previous dialogue state, the current turn dialogue, and the schema graph by internal and external attention mechanisms.
@@
The policy learning module in task-oriented dialogue systems is to generate the following available system action given the state generation result [cuayahuitl2015strategic]. Traditional rule-based methods are first applied in the policy learning procedure [cuayahuitl2015strategic]. Supervised and reinforcement learning are also proven effective in policy learning [su2016continuously, yan2017building]. su2016continuously proposed a two-stage framework for policy learning, i.e., a supervised learning stage and a reinforcement learning stage. chen2019agentgraph proposed a structured deep reinforcement learning approach for policy learning based on graph neural networks. The dialogue policy can be further trained in an end-to-end way with reinforcement learning to lead the system in making policies toward the final performance [yan2017building, chen2019agentgraph]. In an e-commerce scenario, the policy learning component needs to trigger the “recommendation” or a concrete service provided by the customer service [sun2016conversational, sun2018conversational, zhao2021jddc]. Most existing task-oriented dialogue datasets, including WoZ and MultiWoZ, focus mainly on language understanding and dialogue state tracking. However, selecting actions in real life requires obeying user requests and following practical policy limitations. Accordingly, chen2021action presented the action-based conversations dataset consisting of $10042$ conversations containing numerous actions with precise procedural requirements. he2022galaxy utilized semi-supervised pre-training to model explicit dialog policy in task-oriented dialogue systems.
@@
The natural language generation (NLG) component transfers a dialogue action into a natural language response [Chen2017A]. Neural network-based NLG approaches have been proposed for task-oriented dialogues [wen2015stochastic, wen2015semantically, tran2017semantic, zhou2016context]. wen2015stochastic applied an RNN-based generator module and a CNN-based module to rerank candidate utterances. wen2015semantically used an additional control cell to gate the dialogue act to address the slot information omitting and duplicating problems in surface realization. tran2017semantic extended this approach by gating the input token vector of LSTM with the dialogue act. A sequence-to-sequence approach was applied to produce natural language output and deep syntax dependency trees from input dialogue acts [duvsek2016sequence]. Then, zhou2016context proposed an encoder-decoder LSTM-based method to jointly incorporate the request information, semantic slot values, and dialogue act type to generate correct answers. Note that the copy mechanism [vinyals2015pointer, gu2016incorporating] has been successfully applied to the task-oriented dialogue systems to enhance the performance of NLG [eric2017copy, lei2018sequicity, jin2018explicit]. Aiming to augment existing dialog data sets through paraphrase, gao2020paraphrase jointly optimized dialog paraphrase and dialog response generation via a paraphrase augmented response generation approach. Pre-trained language models have shown supreme performances in text generation tasks [li2021pretrained]. In recent years, more and more studies have applied pre-trained language models to enhance the performances of NLG in task-oriented dialogue systems [zhang2019dialogpt, peng-etal-2020-shot, wu2020tod, NEURIPS2020e9462095, zhong2022dialoglm].
@@
End-to-end methods have been proposed for task-oriented dialogue systems. wen2017network proposed an end-to-end trainable goal-oriented dialogue system with a new way of collecting dialogue data based on a pipeline framework toward end-to-end learning for DST and policy learning [zhao2016towards]. The pipeline-aware method can also be implemented and trained end-to-end using the copy mechanism [lei2018sequicity, jin2018explicit, liao2021multi]. A copy-augmented sequence-to-sequence architecture was proposed to provide better performance in task-oriented dialogues [eric2017copy], while eric2017key proposed a key-value retrieval network for task-oriented dialogue response generation. Using the copy mechanism, lei2018sequicity proposed a theoretically and aesthetically framework that is end-to-end trainable using only one seq-to-seq model. jin2018explicit proposed a semi-supervised copy flow neural network to train the end-to-end dialogue generation model. madotto2018mem2seq proposed a memory-to-sequence neural network that combines the multi-hop attention over memories with the idea of a pointer network. xu2018end also applied a pointer network to handle unknown slot values in the absence of a predefined ontology. NEURIPS2020e9462095 enabled modeling of the inherent dependencies between the sub-tasks of task-oriented dialogue by optimizing for all tasks in an end-to-end manner, recasting task-oriented dialogue as a simple and casual language modeling task. liao2021multi proposed a recursive inference mechanism to resolve multi-domain DST in an end-to-end way. Recently, pre-trained language models have also been applied to end-to-end solutions for task-oriented dialogue systems [wu2020tod, lin2021knowledge, 10.1145/3477495.3532069].
@@
In contrast with other types of dialogue systems, evaluation metrics in task-oriented dialogue systems need to consider specific metrics in the experiments: Entity match rate evaluates task completion. According to wen2017network, it determines if a system can generate all correct constraints to search the indicated user entities of the user. This metric is either 0 or 1 for each dialogue. The original success rate metric measures if the system answered all the requested information (e.g., address, phone number) [wen2015stochastic, mrkvsic2015multi]. However, this metric only evaluates recall. As a variant, Success F1 evaluates task completion and is modified from the success rate by balancing both recall and precision [lei2018sequicity]. Automatic user satisfaction received attention in task-oriented dialogues. User simulation is a promising approach to evaluating dialogue systems at scale in task-oriented dialogue scenarios [zhang2020evaluating]. sunzhang2021simulator formulate the task of simulating user satisfaction for evaluating task-oriented dialogue systems to enhance the evaluation of dialogue systems. The authors also shared a dataset about user satisfaction simulation. kim2022mismatch proposed the relative slot accuracy metric in DST evaluation, which is not affected by unseen slots in the current dialogue turn.

### v3

@@ 7.2.1 Introduction to dialogue systems
Generation-based dialogue systems generate natural-sounding replies automatically to exchange information (e.g., knowledge, sentiments, etc.) and complete a variety of specific tasks in a conversational interaction process [Young2013POMDP, shawar2007]. End-to-end textual generation models [shang, vinyals2015neural, sordoni2015, li2016a, li2016b, serban2016building] have proved capable in multiple dialogue systems applications with promising performance. Most end-to-end neural generation models apply an encoder-decoder architecture based on a recurrent neural network, which directly maps an input context to the output response. Several approaches have been proposed to softly model language patterns, such as word alignment and repeating into sequence-to-sequence structure [bahdanau2015neural, gu2016incorporating, serban2017multiresolution, cao2017]. gu2016incorporating propose a copy mechanism to consider additional copying probabilities for contextual words in forum conversations. serban2017multiresolution decode coarse tokens before generating the complete response. Variational neural networks perform efficient inference and learning in models with directed probabilities on a large-scale dataset [Kingma2014Auto, Kingma2014Adam]. cao2017 tackle the boring output issue of deterministic dialogue models by introducing a latent variable model for a one-shot dialogue response. hvred propose HVRED to use the latent variable at the sub-sequence level in a hierarchical setting, whereas ziwwsdm1 add a hierarchical structure and a variational memory module into a neural encoder-decoder network.
@@
Figure 7.5: An example e-commerce dialogue in the JDDC 1.0 dataset. Image source: [chen2020jddc].
@@
In e-commerce platforms, after-sale customer service is the main application scenario for dialogue systems. E-commerce dialogues need to address three targets: (i) task completion, such as changing the order address, providing the receipt, and returning the order; (ii) knowledge-based response selection and generation, such as checking the status of the delivery, answering the request about the refund period; and (iii) empathetic response generation, such as satisfying the consumers’ request and replying to consumers’ complaints. The JDDC datasets have been collected from JD.com, one of the largest e-commerce platforms in China. Fig. 7.5 shows a typical example dialogue from JDDC 1.0 dataset. The blue text shows the target completion task, the red text indicates the knowledge-based response generation task, and the purple text shows the empathetic response generation task. Another characteristic of e-commerce dialogue systems is the phenomenon of multiple modalities. Text and images are often used in customer service dialogues [zhao2021jddc, yuan2022mcic].
@@
Fig. 7.6 shows another example from the JDDC 2.0 dataset to demonstrate the multi-modality characteristic of e-commerce. In this figure, three dialogue segments show images that are used to distinguish different product models for the same brand or used for identifying the location and cause of product failures. According to the above three targets in e-commerce dialogues, we detail e-commerce dialogue systems from three angles: task-oriented dialogue systems, knowledge-grounded dialogue systems, and empathetic dialogue systems in Sections 7.2.27.2.3, and 7.2.4, respectively.
@@ 7.2.2 Task-oriented dialogue systems
Methods underlying task-oriented dialogue systems can be divided into pipeline methods and end-to-end methods. As shown in Fig. 7.4, the pipeline of task-oriented dialogue systems can be divided into natural language understanding (NLU), dialogue state tracking (DST), and dialogue policy learning (DPL), and natural language generation (NLG) [Chen2017A]. For each stage, a number of pipeline-based approaches have been proposed, even though a lot of domain-specific handcrafting in traditional task-oriented dialogue systems is required, which, moreover, is difficult to adapt to new domains [bordes2016learning]. Recently, end-to-end neural network solutions have been widely applied to the task [bordes2016learning, zhao2016towards, wen2017network, jin2018explicit, gao2018neural, madotto2018mem2seq].
@@
In NLU, the dialogue system maps the input utterance into semantic slots. These semantic slots are pre-defined in different application scenarios. NLU includes two challenging problems: intent detection and slot filling. Intent detection methods for language understanding are performed to detect the user’s intent [deng2012use, tur2012towards]. Deep neural networks have also been applied to detect the user’s intent. huang2013learning use convolutional neural networks (CNN) to detect the user’s intent; see also [shen2014learning]. Slot filling is usually set as a sequence labeling problem, where words are assigned with semantic labels [Chen2017A]. Deep belief networks have been successfully applied to address the filling problem [deng2012use, deoras2013deep]. Subsequently, recurrent neural networks have been shown to be effective in slot filling [mesnil2013investigation, sarikaya2011deep, yao2013recurrent, yao2014spoken]. Unlike other NLU approaches, liang2020moss jointly formulate intent detection and slot filling as a sequence generation problem. rastogi2020towards provide a schema-guided paradigm for NLU. Recently, pre-trained language models have been applied to enhance NLU in task-oriented dialogue systems. wu2020tod propose a self-supervised language model trained on multiple task-oriented dialogue system benchmarks. zhang2021effectiveness design a pre-trained model for few-shot NLU by fine-tuning BERT on a small set of labeled utterances. 10.1145/3477495.3532069 explore tree-induced semi-supervised contrastive pre-training for NLU in task-oriented dialogue systems. The authors improve the NLU performance by injecting structural-semantic information to enhance the representation of dialogues. To explore more knowledge from long sequences in dialogue context, zhong2022dialoglm formulate a window-based pre-trained model for NLU based on the sequence-to-sequence model architecture.
@@
In a conversation, a dialogue state refers to a full and temporal representation of each participant’s intention [goddeau1996form]. In task-oriented dialogue generation, dynamically tracking dialogue states is the key to generating coherent and context-sensitive responses. In DST, we use a dialogue state $H_{t}$ to denote the representation of the dialogue till time $t$. Traditional approaches to DST focus on searching hand-crafted rules to select the most likely results [wang2013simple, young2010hidden, williams2012belief], where the dialogue state tracking is transferred to a slot filling problem. This type of slot filling problem has also been addressed by approaches using conditional random fields [lee2013structured, lee2013recipe, ren2013dialog] and maximum entropy [williams2013multi]. However, relying on the most likely results from an NLU module [Perez17], these rule-based systems hardly model uncertainty, which is prone to frequent errors [williams2014web, Perez17].
@@
Unlike rule-based state tracking methods, young2010hidden propose a distributional dialogue state for statistical dialog systems and maintain a distribution over multiple hypotheses facing noisy conditions and ambiguity. Neural networks have been successfully applied to dialogue state tracking [henderson2013deep, mrkvsic2015multi, mrkvsic2017neural]. In task-oriented dialogue systems, end-to-end neural networks are employed for tracking dialogue states via interacting with an external knowledge base [wen2017network, eric2017key, bordes2016learning, williamszweig2017Long]. wen2017network divide the training procedure into two phases: dialogue state tracker training and complete model training. mrkvsic2017neural propose a dialogue state tracker based on word embedding similarities. eric2017key implicitly model a dialogue state through an attention-based retrieval mechanism to reason over a key-value representation of the underlying knowledge base. bordes2016learning track the dialogue context in a memory module and repeatedly search this context to select an adequate system response. Instead of employing symbolic knowledge queries, dhingra-EtAl:2017:Long1 propose an induced “soft” posterior distribution over the knowledge base to search for matching entities. williamszweig2017Long combine an RNN with domain-specific knowledge encoded as software and system action templates. The copying mechanism is shown to be effective as generative dialogue state tracking. lei2018sequicity propose an extendable framework to track dialogue states with a text span, including the constraints for a knowledge base query. The limited amount of labeled data is a severe challenge for DST. jin2018explicit introduce a semi-supervised way to integrate a copy procedure with the dialogue state tracking.
@@
While early studies on DST methods focused on a single-domain scenario, more recent studies have turned their attention to multi-domain DST with the release of a multi-domain DST benchmark dataset, MultiWoZ [budzianowski2018multiwoz]. ramadan2018large jointly identify the domain and tracks the belief states corresponding to that domain to address the multi-domain DST problem. zhou2019multi formulate multi-domain DST as a question-answering task and used reading comprehension techniques to generate the answers. Similarly, gao2019dialog also formulate DST as a reading comprehension task and propose an attention-based neural network to find the state answer as a span over tokens. The DSTC challenges have provided a series of popular experimentation frameworks and dialogue datasets collected through human-machine interactions for benchmarking [henderson2014second, henderson2014third, williams2014dialog, williams2016dialog].
@@
In real-world scenarios, it is often not practical to enumerate all possible slot value pairs and perform scoring from a large, dynamically changing knowledge base [xu2018end]. wu2019transferable propose a method to generate dialogue states from utterances using a copy mechanism, where tracking knowledge across domains is shared. To alleviate data sparsity in DST, yin2020dialog propose a reinforced data augmentation framework to increase both the amount and diversity of the training data. chen2020schema incorporate slot relations and model slot interactions in multi-domain dialogue state tracking to enhance the slot interrelation between disciplines. feng2022dynamic extend this method by dynamically updating slot relations in the schema graph. heck2020trippy maintain two memories in DST: one for system inform slots and one for the previously seen slots. li-etal-2021-generation combine a generation and extraction method with hierarchical ontology integration for DST. To tackle the understanding of ellipsis and reference expressions in open vocabulary-based methods, ouyang2020dialogue propose a copy-augmented encoder-decoder model by connecting the target slot and its source slot explicitly. liao2021multi formulate multi-domain DST as a recursive inference mechanism to improve the generation performance. Most DST models are trained offline, which requires a fixed dataset prepared in advance. Given a new domain in multi-domain DST, campagna2020zero propose a zero-shot transfer learning method to handle new domains without incurring the high cost of data acquisition. More recently, the granularity of dialogue history has been proposed to mitigate the sparseness in DST [yang2021comprehensive]. guo2022beyond propose a multi-perspective dialogue collaborative selector module to dynamically select the granularity of dialogue history in DST.
@@
Pre-trained language models have been shown to be effective in dialogue state tracking. lin2021zero apply a pre-trained model for DST to exploit external knowledge from reading comprehension data. Similarly, zhong2022dialoglm verify that document summarization can provide helpful signals to improve DST. liu2021domain introduce domain-lifelong learning into DST. The authors propose a knowledge preservation network that includes a multi-prototype enhanced retrospection component and a multi-strategy knowledge distillation component. lin2021leveraging successfully apply T5 to improve zero-shot cross-domain DST. lee2021dialogue introduce a solution for multi-domain DST by prompting knowledge from a large-scale pre-trained language model. lin2021knowledge detail a hybrid method to integrate GPT-2 with graph attention networks to enhance the DST performance. To mitigate the problem of incorrect in DST, wang2022luna design a BERT-based method by explicitly aligning each slot with its most relevant utterance.
@@
Scalability, robustness, and efficiency in DST have also been addressed recently. lei2018sequicity formulate a two-stage copy-aware network demonstrating good scalability. ren2019scalable consider the DST task as a sequence generation problem and design a scalable hierarchical encoder-decoder neural network with constant inference time complexity. kumar2020ma extend this to improve the encoding of dialogue context and slot semantics for DST to robustly capture critical dependencies between slots and the conversation history. kim2020efficient focus on an open vocabulary-based setting and consider the dialogue state as a memory that can be selectively overwritten to improve the efficiency in multi-domain DST. zhu2020efficient introduce an efficient multi-domain dialogue state tracker by jointly encoding the previous dialogue state, the current turn dialogue, and the schema graph by internal and external attention mechanisms.
@@
The policy learning module in task-oriented dialogue systems is meant to generate the following available system action given the state generation result [cuayahuitl2015strategic]. Traditional rule-based methods are first applied in the policy learning procedure [cuayahuitl2015strategic]. Supervised and reinforcement learning have also proven to be effective in policy learning [su2016continuously, yan2017building]. su2016continuously propose a two-stage framework for policy learning, i.e., a supervised learning stage and a reinforcement learning stage. chen2019agentgraph propose a structured deep reinforcement learning approach for policy learning based on graph neural networks. The dialogue policy can be further trained in an end-to-end way with reinforcement learning to lead the system in making policies toward the final performance [yan2017building, chen2019agentgraph]. In an e-commerce scenario, the policy learning component needs to trigger the “recommendation” or a concrete service provided by the customer service [sun2016conversational, sun2018conversational, zhao2021jddc]. Most task-oriented dialogue datasets, including WoZ and MultiWoZ, focus on language understanding and dialogue state tracking. However, selecting actions in real life requires obeying user requests and following practical policy limitations. Accordingly, chen2021action present an action-based conversations dataset consisting of $10042$ conversations containing numerous actions with precise procedural requirements. he2022galaxy utilized semi-supervised pre-training to model explicit dialogue policy in task-oriented dialogue systems.
@@
The natural language generation (NLG) component transfers a dialogue action into a natural language response [Chen2017A]. Neural network-based NLG approaches have been proposed for task-oriented dialogues [wen2015stochastic, wen2015semantically, tran2017semantic, zhou2016context]. wen2015stochastic apply an RNN-based generator module and a CNN-based module to rerank candidate utterances. wen2015semantically use an additional control cell to gate the dialogue act to address the slot information omitting and duplicating problems in surface realization. tran2017semantic extend this approach by gating the input token vector of an LSTM with the dialogue act. A sequence-to-sequence approach is applied to produce natural language output and deep syntax dependency trees from input dialogue acts [duvsek2016sequence]. zhou2016context propose an encoder-decoder LSTM-based method to jointly incorporate the request information, semantic slot values, and dialogue act type to generate correct answers. The copy mechanism [vinyals2015pointer, gu2016incorporating] has been successfully applied to the task-oriented dialogue systems to enhance the performance of NLG [eric2017copy, lei2018sequicity, jin2018explicit]. Aiming to augment dialogue datasets through paraphrasing, gao2020paraphrase jointly optimize dialogue paraphrasing and dialogue response generation via a paraphrase augmented response generation approach. Pre-trained language models have shown supreme performance in text generation tasks [li2021pretrained]. In recent years, more and more studies have applied pre-trained language models to enhance the performance of NLG in task-oriented dialogue systems [zhang2019dialogpt, peng-etal-2020-shot, wu2020tod, NEURIPS2020e9462095, zhong2022dialoglm].
@@
End-to-end methods have been proposed for task-oriented dialogue systems. wen2017network propose an end-to-end trainable goal-oriented dialogue system with a new way of collecting dialogue data based on a pipeline framework toward end-to-end learning for DST and policy learning [zhao2016towards]. The pipeline-aware method can also be implemented and trained end-to-end using the copy mechanism [lei2018sequicity, jin2018explicit, liao2021multi]. A copy-augmented sequence-to-sequence architecture has been proposed to provide better performance in task-oriented dialogues [eric2017copy], while eric2017key propose a key-value retrieval network for task-oriented dialogue response generation. Using the copy mechanism, lei2018sequicity formulate a theoretical framework that is end-to-end trainable using only one sequence-to-sequence model. jin2018explicit propose a semi-supervised copy flow neural network to train the end-to-end dialogue generation model. madotto2018mem2seq formulate a memory-to-sequence neural network that combines the multi-hop attention over memories with the idea of a pointer network. xu2018end also apply a pointer network to handle unknown slot values in the absence of a predefined ontology. NEURIPS2020e9462095 enable modeling of the inherent dependencies between the sub-tasks of task-oriented dialogue by optimizing for all tasks in an end-to-end manner, recasting task-oriented dialogues as a simple and casual language modeling task. liao2021multi propose a recursive inference mechanism to resolve multi-domain DST in an end-to-end way. More recently, pre-trained language models have also been applied to end-to-end solutions for task-oriented dialogue systems [wu2020tod, lin2021knowledge, 10.1145/3477495.3532069].
@@
In contrast to other types of dialogue systems, evaluation metrics in task-oriented dialogue systems need to consider specific metrics. Entity match rate evaluates task completion [wen2017network]; it determines if a system can generate all correct constraints to search the indicated entities of the user. This metric is either 0 or 1 for each dialogue. The original success rate metric measures if the system answered all the requested information (e.g., address, phone number) [wen2015stochastic, mrkvsic2015multi]. However, this metric only evaluates recall. As a variant, Success F1 evaluates task completion and is modified from the success rate by balancing both recall and precision [lei2018sequicity]. Automatic user satisfaction has received much attention in task-oriented dialogues. User simulation is a promising approach to evaluate dialogue systems at scale in task-oriented dialogue scenarios [zhang2020evaluating]. sunzhang2021simulator formulate the task of simulating user satisfaction for evaluating task-oriented dialogue systems to enhance the evaluation of dialogue systems. The authors also share a dataset about user satisfaction simulation. kim2022mismatch propose the relative slot accuracy metric in DST evaluation, which is not affected by unseen slots in the current dialogue turn.